# Gubkin str. 8, 119991 Moscow, Russia

### v1

@@ 5.1 Cross-gradients turned off $\mathbf{Y}=0$
$\lambda_{1}^{*}=\frac{8\pi^{2}}{3}\frac{\varepsilon}{J_{11}}\left(1+\sqrt{1-\frac{J_{11}J_{22}}{J_{12}^{2}}}\right),$ (91a)
$\lambda_{2}^{*}=\frac{8\pi^{2}}{3}\frac{\varepsilon}{J_{22}}\left(1-\sqrt{1-\frac{J_{11}J_{22}}{J_{12}^{2}}}\right),$ (91b)
$\lambda_{3}^{*}=\frac{8\pi^{2}}{3}\frac{\varepsilon}{J_{12}},$ (91c)
@@
Turning on $\lambda_{4,5}$ always admits the trivial solution $\lambda_{4}^{\star}=\lambda_{5}^{\star}=0$, which reduces to the parity-symmetric sector. To obtain a genuinely parity-breaking fixed point with $\lambda_{4}^{\star}\neq 0$ and $\lambda_{5}^{\star}\neq 0$, one must impose both conditions $\beta_{\lambda_{4}}=0$ and $\beta_{\lambda_{5}}=0$. Dividing these equations by $\lambda_{4}$ and $\lambda_{5}$ yields a homogeneous linear system for $(\lambda_{4},\lambda_{5})$, which has a nontrivial solution iff its determinant vanishes, i.e. $\Delta=0$ below.
At one loop this produces a one-parameter family (a fixed line), reflecting a marginal direction in the quartic sector. Beyond one loop this degeneracy is expected to be lifted. In the other words for $\lambda_{3}$ satisfying $\Delta_{3}\geq 0$ with $\lambda_{4}^{2},\lambda_{5}^{2}\geq 0$ one can find $\lambda_{1,2,4,5}$ such that $\beta_{\lambda_{i}}=0$. Introducing the notation
@@
$\Delta\equiv\mathcal{X}\mathcal{Y}-J_{11}J_{22}\lambda_{3}^{2},,\,\,\,\,c\equiv\frac{16\pi^{2}}{3}\varepsilon$ (94)
Thus we see that a fully parity-breaking fixed point (with $\lambda_{4},\lambda_{5}\neq 0$ ) must lie on the codimension-1 surface $\mathcal{X}\mathcal{Y}=J_{11}J_{22}\lambda_{3}^{2}$ in $\left(\lambda_{1},\lambda_{2},\lambda_{3}\right)$-space. From $\beta_{\lambda_{1}}=\beta_{\lambda_{2}}=0$ we get the relation between different couplings at fixed point
@@
which also immediately implies constraints
@@
Finally using $\beta_{\lambda_{3}}=0$ and $\Delta=0$ one can obtain different branches of solution for $\lambda_{1,2}$ as functions of $\lambda_{3}$.
Picking a real parameter $\lambda_{3}\neq 0$ for $\sigma=\pm 1$ we define
$\begin{gathered}\Delta_{3}\left(\lambda_{3}\right)\equiv\left(c-2J_{12}\lambda_{3}\right)^{2}-4J_{11}J_{22}\lambda_{3}^{2}\geq 0\\ \mathcal{X}_{\sigma}=\frac{-\left(c-2J_{12}\lambda_{3}\right)+\sigma\sqrt{\Delta_{3}\left(\lambda_{3}\right)}}{2},\quad\mathcal{Y}_{\sigma}=\frac{-\left(c-2J_{12}\lambda_{3}\right)-\sigma\sqrt{\Delta_{3}\left(\lambda_{3}\right)}}{2}\end{gathered}$
@@
The second branch of formal solution has the form and picking $\lambda_{3}\neq 0$ and $\sigma=\pm 1$ let us set
$\mathcal{X}=\lambda_{3}\left(J_{12}+\sigma\sqrt{J_{12}^{2}-J_{11}J_{22}}\right),\quad\mathcal{Y}=\lambda_{3}\left(J_{12}-\sigma\sqrt{J_{12}^{2}-J_{11}J_{22}}\right),$
so $\mathcal{X}\mathcal{Y}=J_{11}J_{22}\lambda_{3}^{2}$ and $\mathcal{X}+\mathcal{Y}=2J_{12}\lambda_{3}$. Then
$\lambda_{1}=\frac{\mathcal{X}+c-2J_{12}\lambda_{3}}{J_{11}},\quad\lambda_{2}=\frac{\mathcal{Y}+c-2J_{12}\lambda_{3}}{J_{22}}.$
and this branch is also compex-valued due to (84).
@@
Also inequalities (95) imposes for $\lambda_{1},\lambda_{2}$ to lie in finite intervals for any given $\lambda_{3}$. So if such fixed points exist, they live in a bounded region of coupling space (at order $\varepsilon$ ). The analysis of stability for these branches repeat the same line of reasoning as for isotropic case and is not affected significantly by anistropy. In the other words the line of fixed points corresponding to the fully interacting $\lambda_{i}$ set of couplings in general has the same stability properties as the isotropic one, changing only location of lines. Anisotropy only restricts the existence of fixed points and shifts their location in coupling space. In particular, for generic initial conditions the RG flow runs away from the fully coupled fixed line.
@@
At one loop in MS the beta functions remain linear in $h$ ( in the $\mathbf{Z}_{t}=\mathbf{1}$, $\mathbf{Y}=0$ limit, the explicit formulas for $\beta_{h_{1}},\dots,\beta_{h_{4}}$ are given above). Therefore the hyperplane $h_{ijk}=0$ is RG-invariant, but it is not IR-attractive: generic infinitesimal cubic perturbations grow toward the IR and destabilize the quartic critical manifolds unless $h$ is tuned to zero (or forbidden by a $\mathbb{Z}_{2}$ symmetry).
@@
and $u_{i}\equiv g_{i}/c$ with $c\equiv\frac{16\pi^{2}}{3}\varepsilon$. Next, rescale the cubics by the $\mu$-independent transformation (App. 270)
@@
independent of the fixed-line parameter $r$ and also independent of the anisotropy ratio $\rho$ (App. C). The corresponding RG exponents are therefore
@@
With the convention $\beta=\mu\frac{d}{d\mu}$, the IR corresponds to $\mu\to 0$, so a perturbation scales as $\delta h\sim\mu^{\omega_{h}}$ and thus grows in the IR when $\omega_{h}<0$. Hence all cubic couplings are relevant perturbations of the quartic fixed line for $\varepsilon>0$. In particular, within the perturbative $4-\varepsilon$ expansion there is no interacting fixed point with $h_{i}^{\star}\neq 0$: at one loop the only fixed locus in the cubic sector is $h^{\star}=0$.
@@
The same conclusion is robust around the generalized Wilson–Fisher fixed point in the parity-symmetric quartic sector: since $\lambda^{\star}=\mathcal{O}(\varepsilon)$, the loop mixing in $\beta_{h}$ is $\mathcal{O}(\varepsilon)\,h$ and cannot overcome the order-one canonical term $(1+\varepsilon/2)h$ for small $\varepsilon$, so cubic perturbations remain relevant near $D=4$.
@@ 5.2 Turning on cross-gradients: $\mathbf{Z}_{t}=\mathbf{1}$ and $\mathbf{Y}\neq 0$
$\mathcal{B}_{ij;kl}\;\equiv\;\varepsilon\,\mathbb{B}_{ij;kl}=\frac{1}{4\pi^{2}}\Big\langle\sum_{a,b\in\{+,-\}}\mathcal{K}\!\big(s_{a}(\hat{\bm{n}}),s_{b}(\hat{\bm{n}})\big)\;\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}\Big\rangle_{S^{2}},$ (115)
@@
(i) Bubble pole as a sum over intermediate two–particle states. At zero external momentum the bubble master integral is
$\mathbb{B}_{ij;kl}(0)=\int\!\frac{d\omega}{2\pi}\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,G_{ik}(\omega,\mathbf{p})\,G_{jl}(\omega,\mathbf{p}),\qquad(D=4-\epsilon,\;d=3-\epsilon),$ (124)
@@
Substituting this into the bubble and integrating over $\omega$ gives the standard two–particle energy denominator:
@@
(ii) The universal kernel $\mathcal{K}$ as a two–particle phase–space density per RG decade. Introduce the positive kernel
$\mathcal{K}\!\big(s_{a},s_{b}\big)\equiv\frac{1}{\sqrt{s_{a}}\sqrt{s_{b}}\big(\sqrt{s_{a}}+\sqrt{s_{b}}\big)}=\frac{1}{v_{a}v_{b}(v_{a}+v_{b})}.$ (130)
Using $E_{a}=|\mathbf{p}|v_{a}$ in (129) one finds
$\frac{1}{2E_{a}E_{b}(E_{a}+E_{b})}=\frac{1}{2|\mathbf{p}|^{3}}\,\mathcal{K}(s_{a},s_{b}).$ (131)
Thus the radial integral in $d=3$ becomes $\int d^{3}\mathbf{p}\,|\mathbf{p}|^{-3}\sim\int dp\,p^{-1}$, so the UV divergence is logarithmic: in a Wilsonian cutoff scheme it is $\ln b$ from integrating out a thin momentum shell $\Lambda/b<|\mathbf{p}|<\Lambda$, while in DR+MS it appears as $1/\epsilon$.
A particularly illuminating representation rewrites the energy denominator $1/(E_{a}+E_{b})$ as an integral over the total two–particle energy $\Omega$:
$\frac{1}{E_{a}+E_{b}}=\int_{0}^{\infty}\frac{d\Omega}{\Omega}\,\delta\!\big(\Omega-(E_{a}+E_{b})\big).$ (132)
Then
$\frac{1}{2E_{a}E_{b}(E_{a}+E_{b})}=2\int_{0}^{\infty}\frac{d\Omega}{\Omega}\,\frac{1}{4E_{a}E_{b}}\,\delta\!\big(\Omega-(E_{a}+E_{b})\big),$ (133)
and the bubble becomes
$\mathbb{B}_{ij;kl}(0)=2\int_{0}^{\infty}\frac{d\Omega}{\Omega}\,\sum_{a,b=\pm}\,\Phi^{(2)}_{ab;\,ij;kl}(\Omega),$ (134)
where we introduced the (directionally resolved) two–particle phase–space density
$\Phi^{(2)}_{ab;\,ij;kl}(\Omega)\equiv\int\!\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,\frac{\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}}{4E_{a}(\mathbf{p})E_{b}(\mathbf{p})}\,\delta\!\big(\Omega-(E_{a}(\mathbf{p})+E_{b}(\mathbf{p}))\big).$ (135)
Equation (134) is a Wilsonian statement: the one–loop correction is an integral over all intermediate two–particle energies $\Omega$, weighted by $d\Omega/\Omega$ (one RG decade) times the density of intermediate states.
For our linear dispersions in $d=3$ the phase–space density at fixed $\Omega$ is scale–invariant: evaluating (135) in spherical variables $\mathbf{p}=\rho\,\hat{\bm{n}}$ gives
$\Phi^{(2)}_{ab;\,ij;kl}(\Omega)$ $=\int\!\frac{d\Omega_{2}}{(2\pi)^{3}}\int_{0}^{\infty}d\rho\,\frac{\rho^{2}}{4(\rho v_{a})(\rho v_{b})}\,\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}\,\delta\!\big(\Omega-\rho(v_{a}+v_{b})\big)$
$=\int\!\frac{d\Omega_{2}}{(2\pi)^{3}}\,\frac{\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}}{4v_{a}v_{b}(v_{a}+v_{b})}=\frac{1}{8\pi^{2}}\,\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}\,\mathcal{K}(s_{a},s_{b}).$
Thus the universal kernel is nothing but the (directional) two–particle phase–space weight:
$\frac{1}{8\pi^{2}}\,\mathcal{K}(s_{a},s_{b})=\int\!\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,\frac{1}{4E_{a}E_{b}}\,\delta\!\big(\Omega-(E_{a}+E_{b})\big)\qquad(d=3,\;\text{linear }E_{a}).$ (136)
Substituting this back into (134) shows explicitly that the MS pole coefficient $\mathcal{B}=\epsilon\,\mathbb{B}$ is the two–particle phase–space density per RG decade: the $1/\epsilon$ pole in DR corresponds to the logarithmic $\int d\Omega/\Omega$ integral.
(iii) Geometric picture: integration over constant–energy surfaces. The representation (135) may also be rewritten as an integral over the constant–energy surface $\Sigma^{ab}_{\Omega}\equiv\{\mathbf{p}:\ E_{a}(\mathbf{p})+E_{b}(\mathbf{p})=\Omega\}$:
$\int d^{3}\mathbf{p}\,F(\mathbf{p})\,\delta\!\big(\Omega-(E_{a}+E_{b})\big)=\int_{\Sigma^{ab}_{\Omega}}\frac{dS}{|\nabla_{\mathbf{p}}(E_{a}+E_{b})|}\,F(\mathbf{p}).$ (137)
@@
$\Pi_{+}(\hat{\bm{n}})=\begin{pmatrix}\cos^{2}\theta&\sin\theta\cos\theta\\ \sin\theta\cos\theta&\sin^{2}\theta\end{pmatrix},\qquad\Pi_{-}(\hat{\bm{n}})=\begin{pmatrix}\sin^{2}\theta&-\sin\theta\cos\theta\\ -\sin\theta\cos\theta&\cos^{2}\theta\end{pmatrix}$. (139)
Define the “populations” and “coherence” of mode $a$ by
@@
Then $p_{a},q_{a}\in[0,1]$ measure how much of $\phi$ or $\chi$ participates in the fast mode $a$, while $r_{a}=\pm\frac{1}{2}\sin 2\theta$ measures the off–diagonal overlap (coherence). Under $\mathbf{Y}\to-\mathbf{Y}$ one has $\theta\to-\theta$, hence $p_{a},q_{a}$ are even while $r_{a}$ is odd. This parity property will immediately explain which weights are even/odd in $\mathbf{Y}$.
(v) The six weights as phase–space–weighted moments of populations and coherence. In the $Z_{t}=\mathbf{1}$ truncation, the residue $\mathcal{B}$ can be viewed as a symmetric $3\times 3$ matrix acting on the space of symmetric bilinears $(\phi^{2},\chi^{2},\sqrt{2}\,\phi\chi)$. Equivalently, the six independent entries of $\mathcal{B}$ may be packaged into the six scalar weights defined in (116). A convenient explicit form is
@@
(vi) Positivity and useful inequalities. Since $\mathcal{K}(s_{a},s_{b})>0$ and the projectors are real, $\mathcal{B}$ defines a positive bilinear form on symmetric bilinears, hence its $3\times 3$ matrix is positive semidefinite. As a result, the JJ–weights satisfy Cauchy–Schwarz–type inequalities (useful as consistency checks), for example
@@
(vii) How the weights enter $\beta$–functions (diagrammatic meaning). At one loop the quartic beta functions arise from fish diagrams and are quadratic in $\lambda$. At the tensor level
@@
$w_{++}=\mathcal{K}(s_{+},s_{+})=\frac{1}{2s_{+}^{3/2}},\qquad w_{--}=\mathcal{K}(s_{-},s_{-})=\frac{1}{2s_{-}^{3/2}}$ (150)
@@
These identities immediately imply nontrivial relations among the six weights, e.g.

### v2

@@ 5.1 Cross-gradients turned off: $\mathbf{Y}=0$
$\lambda_{1}^{*}=\frac{8\pi^{2}}{3}\frac{\epsilon}{J_{11}}\left(1+\sqrt{1-\frac{J_{11}J_{22}}{J_{12}^{2}}}\right),$ (89a)
$\lambda_{2}^{*}=\frac{8\pi^{2}}{3}\frac{\epsilon}{J_{22}}\left(1-\sqrt{1-\frac{J_{11}J_{22}}{J_{12}^{2}}}\right),$ (89b)
$\lambda_{3}^{*}=\frac{8\pi^{2}}{3}\frac{\epsilon}{J_{12}},$ (89c)
@@
Turning on $\lambda_{4,5}$ always admits the trivial solution $\lambda_{4}^{*}=\lambda_{5}^{*}=0$, which reduces to the parity-symmetric sector. To obtain a genuinely parity-breaking fixed point with $\lambda_{4}^{*}\neq 0$ and $\lambda_{5}^{*}\neq 0$, one must impose both conditions $\beta_{\lambda_{4}}=0$ and $\beta_{\lambda_{5}}=0$. Dividing these equations by $\lambda_{4}$ and $\lambda_{5}$ yields a homogeneous linear system for $(\lambda_{4},\lambda_{5})$, which has a nontrivial solution iff its determinant vanishes, i.e. $\Delta=0$ below.
At one loop this produces a one-parameter family (a fixed line), reflecting a marginal direction in the quartic sector. Beyond one loop this degeneracy is expected to be lifted. In other words, for $\lambda_{3}$ satisfying $\Delta_{3}\geq 0$ (determinant $\Delta_{3}$ will also be defined below) with $\lambda_{4}^{2},\lambda_{5}^{2}\geq 0$ one can find $\lambda_{1,2,4,5}$ such that $\beta_{\lambda_{i}}=0$. Introducing the notation
@@
$\Delta\equiv\mathcal{X}\mathcal{Y}-J_{11}J_{22}\lambda_{3}^{2},\,\,\,\,c\equiv\frac{16\pi^{2}}{3}\epsilon.$ (92)
Thus we see that a fully parity-breaking fixed point (with $\lambda_{4},\lambda_{5}\neq 0$ ) must lie on the codimension-1 surface $\mathcal{X}\mathcal{Y}=J_{11}J_{22}\lambda_{3}^{2}$ in $\left(\lambda_{1},\lambda_{2},\lambda_{3}\right)$-space. From $\beta_{\lambda_{1}}=\beta_{\lambda_{2}}=0$ we get the relations between different couplings at fixed point
@@
which also immediately imply constraints
@@
Finally, using $\beta_{\lambda_{3}}=0$ and $\Delta=0$ one can obtain different branches of solution for $\lambda_{1,2}$ as functions of $\lambda_{3}$.
(i) For the first branch, the equation $\beta_{\lambda_{3}}=0$ together with $\Delta=0$ gives:
$\mathcal{X}+\mathcal{Y}=2J_{12}\lambda_{3}-c,$ (94)
and
$\mathcal{X}\mathcal{Y}=J_{11}J_{22}\lambda_{3}^{2}.$ (95)
Then picking a real parameter $\lambda_{3}\neq 0$ for $\sigma=\pm 1$ we solve these two equations together and also define
$\begin{gathered}\Delta_{3}\left(\lambda_{3}\right)\equiv\left(c-2J_{12}\lambda_{3}\right)^{2}-4J_{11}J_{22}\lambda_{3}^{2}\geq 0,\\ \mathcal{X}_{\sigma}=\frac{-\left(c-2J_{12}\lambda_{3}\right)+\sigma\sqrt{\Delta_{3}\left(\lambda_{3}\right)}}{2},\quad\mathcal{Y}_{\sigma}=\frac{-\left(c-2J_{12}\lambda_{3}\right)-\sigma\sqrt{\Delta_{3}\left(\lambda_{3}\right)}}{2},\end{gathered}$
@@
(ii) The second branch of formal solution has the following form (with $\lambda_{3}\neq 0$ and $\sigma=\pm 1$):
$\mathcal{X}_{\sigma}=\lambda_{3}\left(J_{12}+\sigma\sqrt{J_{12}^{2}-J_{11}J_{22}}\right),\quad\mathcal{Y}_{\sigma}=\lambda_{3}\left(J_{12}-\sigma\sqrt{J_{12}^{2}-J_{11}J_{22}}\right),$
so $\mathcal{X}_{\sigma}\mathcal{Y}_{\sigma}=J_{11}J_{22}\lambda_{3}^{2}$ and $\mathcal{X}_{\sigma}+\mathcal{Y}_{\sigma}=2J_{12}\lambda_{3}$. Then, using (90)
$\lambda_{1}=\frac{\mathcal{X}_{\sigma}+c-2J_{12}\lambda_{3}}{J_{11}},\quad\lambda_{2}=\frac{\mathcal{Y}_{\sigma}+c-2J_{12}\lambda_{3}}{J_{22}},$
and this branch is also complex-valued due to (82).
@@
Also inequalities (93) imposes for $\lambda_{1},\lambda_{2}$ to lie in finite intervals for any given $\lambda_{3}$. So if such fixed points exist, they live in a bounded region of coupling space (at order $\epsilon$). The analysis of stability for these branches repeat the same line of reasoning as for isotropic case and is not affected significantly by anisotropy. In other words, the line of fixed points corresponding to the fully interacting $\lambda_{i}$ set of couplings in general has the same stability properties as the isotropic one, changing only location of lines. Anisotropy only restricts the existence of fixed points and shifts their location in coupling space. In particular, for generic initial conditions the RG flow runs away from the fully coupled fixed line.
@@
At one loop in MS the beta functions remain linear in $h$ ( in the $\mathbf{Z}_{t}=\mathbf{1}$, $\mathbf{Y}=0$ limit, the explicit formulas for $\beta_{h_{1}},\dots,\beta_{h_{4}}$ are given by (84)). Therefore the hyperplane $h_{ijk}=0$ is RG-invariant, but it is not IR-attractive: generic infinitesimal cubic perturbations grow toward the IR and destabilize the quartic critical manifolds unless $h$ is tuned to zero (or forbidden by a $\mathbb{Z}_{2}$ symmetry).
@@
and $u_{i}\equiv g_{i}/c$ with $c\equiv\frac{16\pi^{2}}{3}\varepsilon$. Next, rescale the cubics by the $\mu$-independent transformation (App. 269)
@@
where “spec” means spectrum and it is independent of the fixed-line parameter $r$ and also independent of the anisotropy ratio $\rho$ (App. C). The corresponding RG exponents are therefore
@@
With the convention $\beta_{g}=\mu\frac{dg}{d\mu}$ (with fixed bare coupling), the IR corresponds to $\mu\to 0$, so a perturbation scales as $\delta h\sim\mu^{\omega_{h}}$ and thus grows in the IR when $\omega_{h}<0$. Hence all cubic couplings are relevant perturbations of the quartic fixed line for $\varepsilon>0$. In particular, within the perturbative $4-\varepsilon$ expansion there is no interacting fixed point with $h_{i}^{*}\neq 0$: at one loop the only fixed locus in the cubic sector is $h^{*}=0$.
@@
The same conclusion is robust around the generalized Wilson–Fisher fixed point in the parity-symmetric quartic sector: since $\lambda^{*}=\mathcal{O}(\varepsilon)$, the loop mixing in $\beta_{h}$ is $\mathcal{O}(\varepsilon)\,h$ and cannot overcome the order-one canonical term $(1+\varepsilon/2)h$ for small $\varepsilon$, so cubic perturbations remain relevant near $D=4$.
@@ 5.2 Turning on cross-gradients: $\mathbf{Z}_{t}=\mathbf{1}$ and $\mathbf{Y}\neq 0$
$\mathcal{B}_{ij;kl}\;\equiv\;\varepsilon\,\mathbb{B}_{ij;kl}=\frac{1}{4\pi^{2}}\Big\langle\sum_{a,b\in\{+,-\}}\mathcal{K}\!\big(s_{a}(\hat{\bm{n}}),s_{b}(\hat{\bm{n}})\big)\;\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}\Big\rangle_{S^{2}}.$ (115)
@@
Bubble pole as a sum over intermediate two–particle states. At zero external momentum the bubble master integral is
$\mathbb{B}_{ij;kl}(0)=\int\!\frac{d\omega}{2\pi}\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,G_{ik}(\omega,\mathbf{p})\,G_{jl}(\omega,\mathbf{p}),$ (124)
@@
Substituting this into the bubble (124) and integrating over $\omega$ gives the standard two–particle energy denominator:
@@
Spectral (Källén–Lehmann–type) viewpoint. It is useful to view the bubble (3) as a Euclidean two–point function of a bilinear field density. To make the spectral structure explicit, consider the bubble (3) with a purely Euclidean external energy $k_{0}$ and vanishing external spatial momentum:
$\mathbf{B}_{ij;kl}(k_{0})\;\equiv\;\int\!\frac{d\omega}{2\pi}\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,G_{ik}(\omega,\mathbf{p})\,G_{jl}(\omega+k_{0},\mathbf{p}).$ (130)
In the massless UV regime we may use the normal–mode decomposition
$G(\omega,\mathbf{p})=\sum_{a}\frac{\Pi_{a}(\hat{\bm{n}})}{(\omega^{2}+E_{a}(\mathbf{p})^{2})},$
@@
$\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\frac{1}{(\omega^{2}+E_{a}^{2})\big((\omega+k_{0})^{2}+E_{b}^{2}\big)}=\frac{E_{a}+E_{b}}{2E_{a}E_{b}\big(k_{0}^{2}+(E_{a}+E_{b})^{2}\big)},$ (131)
thus we arrive at
$\mathbf{B}_{ij;kl}\left(k_{0}\right)=\sum_{a,b}\int\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\Pi_{a}(\hat{\mathbf{n}})_{ik}\Pi_{b}(\hat{\mathbf{n}})_{jl}\frac{E_{a}+E_{b}}{2E_{a}E_{b}\left(k_{0}^{2}+\left(E_{a}+E_{b}\right)^{2}\right)}.$ (132)
Next, introduce the positive two–particle spectral density
$\rho^{(2)}_{ij;kl}(\Omega)\;\equiv\;\sum_{a,b}\int\!\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\,\frac{\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}}{4E_{a}(\mathbf{p})E_{b}(\mathbf{p})}\,\delta\!\big(\Omega-(E_{a}(\mathbf{p})+E_{b}(\mathbf{p}))\big),$
$\rho_{ij;kl}^{(2)}(\Omega)\succeq 0,$
and the positivity is immediate, since $\Pi_{a}$ are Hermitian projectors and $E_{a}>0$. Then the bubble (130) admits the following Källén–Lehmann–type representation
$\mathbf{B}_{ij;kl}(k_{0})=\int_{0}^{\infty}d\Omega^{2}\,\frac{\rho^{(2)}_{ij;kl}(\Omega)}{k_{0}^{2}+\Omega^{2}}\,.$ (133)
At $k_{0}=0$ this reduces to $\mathbf{B}_{ij;kl}(0)=2\int_{0}^{\infty}\frac{d\Omega}{\Omega}\,\rho^{(2)}_{ij;kl}(\Omega)$, which makes explicit the logarithmic UV sensitivity ($\int d\Omega/\Omega$) that becomes a $1/\epsilon$ pole in DR+MS.
Evaluating $\rho_{ij;kl}^{(2)}$ in spherical variables $\mathbf{p}=\rho\,\hat{\bm{n}}$ gives an $\Omega$–independent integrand and reproduces the universal kernel:
$\rho^{(2)}_{ij;kl}=\frac{1}{8\pi^{2}}\Big\langle\sum_{a,b}\Pi_{a}(\hat{\bm{n}})_{ik}\,\Pi_{b}(\hat{\bm{n}})_{jl}\,\mathcal{K}\!\big(s_{a}(\hat{\bm{n}}),s_{b}(\hat{\bm{n}})\big)\Big\rangle_{S^{2}},$ (134)
$\mathcal{K}(s_{a},s_{b})=\frac{1}{v_{a}v_{b}(v_{a}+v_{b})}.$
Thus $\mathcal{K}$ controls the spectral weight of intermediate two–particle states, while the projector entries provide the coherence factors (overlaps with the microscopic fields). Moreover, using (133) one may straightforwardly show, that in a Lorentz’s signature continuation $\rho_{ij;kl}^{(2)}$ becomes the discontinuity across the corresponding two–particle cut.
@@
$\Pi_{+}(\hat{\bm{n}})=\begin{pmatrix}\cos^{2}\theta&\sin\theta\cos\theta\\ \sin\theta\cos\theta&\sin^{2}\theta\end{pmatrix},\quad\Pi_{-}(\hat{\bm{n}})=\begin{pmatrix}\sin^{2}\theta&-\sin\theta\cos\theta\\ -\sin\theta\cos\theta&\cos^{2}\theta\end{pmatrix}$. (136)
[…]
@@
Then $p_{a},q_{a}\in[0,1]$ measure how much of $\phi$ or $\chi$ participates in the fast mode $a$, while $r_{a}=\pm\frac{1}{2}\sin 2\theta$ measures the off–diagonal overlap, i.e. coherence. Under $\mathbf{Y}\to-\mathbf{Y}$ one has $\theta\to-\theta$, hence $p_{a},q_{a}$ are even while $r_{a}$ is odd. This parity property will immediately explain which weights are even/odd in $\mathbf{Y}$.
The six weights as phase–space–weighted moments of populations and coherence. In the $Z_{t}=\mathbf{1}$ truncation, the residue $\mathcal{B}$ can be viewed as a symmetric $3\times 3$ matrix acting on the space of symmetric bilinears $(\phi^{2},\chi^{2},\sqrt{2}\,\phi\chi)$. Equivalently, the six independent entries of $\mathcal{B}$ may be packaged into the six scalar weights defined in (116). A convenient explicit form is
@@
Positivity and useful inequalities. Since $\mathcal{K}(s_{a},s_{b})>0$ and the projectors are real, $\mathcal{B}$ defines a positive bilinear form on symmetric bilinears, hence its $3\times 3$ matrix is positive semidefinite. As a result, the JJ–weights satisfy Cauchy–Bunyakovsky–Schwarz–type inequalities (useful as consistency checks), for example
@@
How the weights enter $\beta$–functions (diagrammatic meaning). At one loop the quartic beta functions arise from fish diagrams and are quadratic in $\lambda$. At the tensor level
@@
$w_{++}=\mathcal{K}(s_{+},s_{+})=\frac{1}{2s_{+}^{3/2}},$ (147)
$w_{--}=\mathcal{K}(s_{-},s_{-})=\frac{1}{2s_{-}^{3/2}},$ (148)
$w_{+-}=K(s_{+},s_{-})=\frac{1}{\sqrt{s_{+}}\sqrt{s_{-}}\left(\sqrt{s_{+}}+\sqrt{s_{-}}\right)}.$ (149)
@@
These identities immediately imply nontrivial relations among the six weights: